Determine the pHfor the following solutions:

Part B

[H3O+] = 4.2×10−3 M

Part C

[H3O+]= 2×10−4M

Part D

[OH−]= 8.8×10−9M
Answer #1

Here , for part B & C , the formula needed is ,

For part D , two equations are needed .

Part B

[H3O+]=4.2*10-3 M

[ANSWER]

Part C

[H3O+]=2*10-4 M

[ANSWER]

Part D

[OH-]=8.8*10-9

[ANSWER]